the measure of an angle and its supplement are in a ratio of is 1:8
find the measure of the angle?

Answer by checkley77(12569)

1n+8n=180
9n=180
n=180/9=n=20 degrees ans. for the smaller angle.
180-n=160 degrees ans. for the larger angle.
Proof:
1/8=20/160
160=160